OpenStudy (anonymous):

j(x) = k cos(x), x ≤ 0 11ex − k, 0 < x find K

OpenStudy (anonymous):

\[j(x)= Kcos(x) \] at x<or equal 0

OpenStudy (anonymous):

given that it is continuous ?

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} k\cos(x) & \text{if} & x\leq0 \\ 11e^x-k& \text{if} & x> 0 \end{array} \right.\] ??

OpenStudy (anonymous):

this is easier than you may think replace \(x\) by zero in both expressions, set them equal and solve for \(k\)